How far does a sled travel in 5 seconds while accelerating from 4m/s to 10m/s. *


1
Expert's answer
2021-04-29T10:43:03-0400
"s=\frac{v_i+v_f}{2}t\\s=\frac{4+10}{2}5=35\ m"